# Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging

### v2

@@ Abstract
We uncover a new mechanism whereby the triple interplay of non-Hermitian pumping, bosonic interactions and nontrivial band topology leads to ultra-strong bosonic condensation. The extent of condensation goes beyond what is naively expected from the interaction-induced trapping of non-Hermitian pumped states, and is based on an emergent caging mechanism that can be further enhanced by topological boundary modes. Beyond our minimal model with 2 bosons, this caging remains applicable for generic many-boson systems subject to a broad range of density interactions and non-Hermitian hopping asymmetry. Our novel new mechanism for particle localization and condensation would inspire fundamental shifts in our comprehension of many-body non-Hermitian dynamics and opens new avenues for controlling and manipulating bosons.
Introduction.— There has been growing excitement in interplay between many-body interactions and the non-Hermitian skin effect (NHSE) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. This stems from the non-locality and sensitivity of the NHSE [22, 23, 14, 24, 25, 26, 3, 27, 28, 5], which profoundly impacts particle localization [1, 2, 3, 4, 5, 6, 7] and condensation dynamics [8, 29, 9, 10, 11, 12], which leads to unexpected consequences in the presence of quantum many-body interactions. These include the emergence of a many-fermion Fermi skin [2, 7], exotic spin liquids from environmental couplings [1], spectral symmetry breaking [13], unconventional quantum dynamics in non-Hermitian baths [9, 30, 10, 31, 32, 33] and the non-Hermitian Mott skin effect [20].
In this work, we uncovered a new mechanism whereby bosons can be made to condense very strongly through the triple interplay of non-Hermitian pumping, density interactions and non-trivial band topology. Due to the emergent non-locality of both the NHSE [22, 23, 14, 24, 25] and the density interactions, the observed phenomenon contrasts starkly with the conventional behavior of interacting bosons, such as photons, which repel each other when they interact i.e. photon blockade [34, 35, 36, 37]. In particular, we show that ultra-strong localization, and not just boundary state accumulation, arises from an emergent caging mechanism that requires both non-Hermitian pumping and density interactions.
Ultra-strong non-Hermitian Bosonic condensation.— To demonstrate how non-Hermitian pumping, topology and bosonic interactions can interplay to cause unexpectedly strong bosonic condensation, we consider a minimal interacting 1D bosonic lattice model that contains topological edge modes at its boundary sites:
@@
which is a Su-Schrieffer-Heeger (SSH) lattice [38] with asymmetric intercell ($t_{L}$ and $t_{R}$) hoppings and unit intracell hoppings, equipped with a density interaction $\frac{1}{2}Un_{x_{0}}^{2}$ at site $x_{0}$, $U\geq 0$. Here $b_{x}$ ($b_{x}^{\dagger}$) is the bosonic annihilation (creation) operator at sites $x=1,2,\ldots,2L$, with $n_{x_{0}}=b_{x_{0}}^{\dagger}b_{x_{0}}$ the boson number operator. Non-Hermitian pumping [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 24, 50, 51, 52, 53, 54, 22, 55, 56, 57, 58, 59, 60, 61, 14, 62, 63, 64, 65, 66, 67, 68, 69, 70] occurs towards the left if $r=t_{L}/t_{R}>1$, accumulating the bosons towards the left boundary. However, when there are multiple indistinguishable bosons, this accumulation would also be affected by the repulsive ($U>0$) interaction at site $x_{0}$, with strength depending on bosonic occupancy. In this work, we have chosen the interaction $\frac{1}{2}Un_{x_{0}}^{2}$ that mimics the simplest possible nonlinearity in the mean-field limit, different from the usual Bose-Hubbard interaction [71, 34, 72, 35, 36, 37] by a local density shift [73, 74, 75, 76, 77].
Figure 1: Spatial density accumulation (a-d) and ultra-strong bosonic condensation (e) due to a boundary density interaction at $x_{0}=1$, for different two-boson initial states (1,1), (4,4), and (1,4) evolving via $H$ [Eq. 1]. $r=1$ or $4$ corresponds to Hermitian (a,b) or non-Hermitian (c,d) cases, and $U=0$ or $U\neq 0$ corresponds the absence (a,c) or presence (b,d) of the density interaction. (a-d) Heatmaps depict $\rho_{x}(t)$ [Eq. 3], the time-dependent boson number over the simulation duration $t\in[0,T]$ with $T=10$. The line plot above each heatmap shows the corresponding time-averaged boson number profile $\bar{\rho}_{i}$ [Eq. 4]. The bosons are most likely to be trapped at site $x=1$ in the presence of both leftwards NHSE ($r>0$) and nonzero interaction ($U>0$), as shown by the white and bright yellow regions in the heatmaps. While normally being considered as repulsive, the interaction term $U$ also acts as a trap at $x_{0}=1$ in non-equilibrium dynamics. (e) The dramatic interplay of NHSE pumping and bosonic interaction is most saliently expressed through the bosonic condensation at $x=1$, as given by the two-boson correlation $P_{11}$ [Eq. 5]. Here, we observe significantly larger $P_{11}$, indicated by the red curves in all three (e) subfigures, which demonstrates the greatly enhanced ability to trap both bosons at site 1 when both $U>0$ and $r>0$. $t_{0}=3$ for all cases. $r=4$ corresponds to $t_{L}=1.6$ and $t_{R}=0.4$.
@@
where $v_{xx^{\prime}}(t)=v_{x^{\prime}x}(t)$ is the amplitude of the basis state $|(x,x^{\prime})\rangle=b_{x}^{\dagger}b_{x^{\prime}}^{\dagger}|0\rangle$ containing indistinguishable bosons at sites $x,x^{\prime}$. We define the 2-boson density at site $x=1,2,\ldots,2L$ by
@@
with $0\leq\rho_{x}(t)\leq 2$. Associated with it is the time averaged spatial density
@@
where $T$ is the simulation duration. While $\rho_{x}(t)$ and $\bar{\rho}_{x}$ reveal where the bosons localize on the lattice, a high value of $\rho_{x}(t)$ or $\bar{\rho}_{x}$ can physically arise either due to strong single-boson localization at site $x$, or moderate double-boson condensation at the same site. To quantify this important distinction, we also examine the two-boson correlation probability
@@
which represents the probability of observing a boson at sites $x$ and $x^{\prime}$ at the same time $t$. For revealing underlying trends in the evolution of the two-boson correlation, we also define the time-smoothed correlation
@@
We first consider scenarios where the density interaction is at the leftmost (boundary) site $x_{0}=1$. Figs. 1(a-d) showcase the evolution of the dynamical density $\rho_{x}(t)$ [Eq. 3] and its time-average $\bar{\rho}_{x}$ [Eq. 4] of three illustrative initial 2-boson state configurations. For the initial state $(1,1)$ (Top Row) where both bosons are already on $x_{0}=1$, the density interaction indeed repels the bosons in the Hermitian limit, as evidenced in the suppressed $\bar{\rho}_{x}$ density at $x=x_{0}=1$ across the Hermitian interacting vs. non-interacting cases (orange vs. blue). Somewhat expectedly, this suppression vanishes when the NHSE counteracts the repulsion by pumping the bosons towards $x_{0}=1$ (green and red). However, for the initial state $(4,4)$ (Center Row) where both bosons are initially far from $x_{0}=1$, the boson density oscillates and spreads out, failing to accumulate appreciably at $x_{0}=1$ except when both the non-Hermiticity and interaction are present (red). Indeed, in this non-equilibrium scenario, the $U>0$ density term behaves more like a local potential well that traps the $\rho_{x}(t)$ at $x=x_{0}$, rather than a repulsion. Qualitatively similar behavior is observed in the density evolution for initial state $(1,4)$ (Bottom Row), when one of the bosons is initially already at $x_{0}=1$ and is as such unaffected by the NHSE.
@@
What is striking, however, is the ultra-strong two-boson condensation at boundary site $x_{0}=1$ when non-Hermiticity ($r>1$) and the density interaction ($U>0$) are both present. This is revealed in the dynamical $P_{11}(t)$ [Eq. 5] plots in Fig. 1(e), which shows the probability of both bosons condensing at $x=1$. From the Top Row, when both bosons are prepared at $x=1$ (initial state =(1,1)), they can only remain both there if the interaction and non-Hermiticity are simultaneously present (red). But most surprising is what happens when at least one boson is initially away from $x=1$ (Center and Bottom Rows): we observe overwhelmingly higher $P_{11}$ only when non-Hermiticity and interaction simultaneously interplay (red), compared to having either on their own (yellow, green). This suggests that trapping both bosons and suppress photon blockade [35, 36, 34, 37] hinges on an emergent consequence of this interplay, a fact not evident from the density evolution alone [Fig. 1(a-d)].
@@
To quantitatively characterize this two-boson condensation, we plot in Fig. 2(a) the time-averaged condensation probability $\bar{P}_{11}=\bar{P}_{11}(T)$ with $\Delta t=T$ [Eq. 6], which is the correlation $P_{11}$ smoothed over the entire simulation duration $T$. The heatmap of $\bar{P}_{11}$ in the interaction strength vs. non-Hermiticity $U$-$r$ plane [Fig. 2(a)] reveals significantly enhanced $\bar{P}_{11}$ (colored) for certain $(U,r)$ regions, compared to the $r=1$ Hermitian limit where $\bar{P}_{11}$ essentially vanishes (black). While $\bar{P}_{11}$ generally increases with $r$, optimal condensation $\bar{P}_{11}$ occurs at windows of $U$ that are highly dependent on the initial state, as will be explained below.
@@
Figure 2: Extent of two-boson condensation due to a boundary density interaction at $x_{0}=1$ [Eq. 1], and its correspondence with the boundary localization of 2-boson spectral bands. (a) Time-averaged two-boson condensation probability $\bar{P}_{11}$ [Eq. 6] at site $x=1$, in the parameter space of non-Hermitian hopping asymmetry $r$ and density interaction strength $U$. (b) Corresponding two-boson spectra at $r=4(t_{L}=1.6$ and $t_{R}=0.4),t_{0}=3$, which features five bands, with bands 2 and 4 being containing a pair of interaction-hybridized bulk and topological bosons. As $U$ increases, it creates a group of eigenstates at $E\approx U$ that exhibits strong localization $\psi_{11}$ (red) at $x=x_{0}=1$. Strong $\psi_{11}$ localization in the hybrid topological band 4 leads to suppressed $\bar{P}_{11}$ for the boundary-localized initial state $(1,1)$ (circled in dashed black), but enhanced $\bar{P}_{11}$ condensation for the initial state $(1,4)$ (solid black). By contrast, strong $\psi_{11}$ localization in the bulk bands 3 and 5 (circled in light blue) corresponds to enhanced $\bar{P}_{11}$ for the bulk initial state $(4,4)$.
Figure 3: Ultrastrong boundary condensation $\bar{P}_{11}$ from interaction-induced caging, for initial state (2,2) within a $x_{0}=3$ cage for a 5 unit cell chain. (a) Snapshots of the time-smoothed ($\Delta t=2$) two-boson correlation probability $\bar{P}_{11}$ [Eq. 6] for (a1) non-interacting and (a2) interacting cases. The L-shaped cage in (a2) traps the bosons tightly. (b1) Evolution of probability of both bosons at the left boundary, with greatly enhanced $\bar{P}_{11}(t)$ only for the non-Hermitian interacting case (red). (b2) When only one boson is present, the boundary density $\bar{\rho}_{1}(t)$ remains low regardless of whether a $U$ barrier is present, showcasing that the caging mechanism is an interaction effect.
Topological origin of ultra-strong bosonic condensation.— Intriguingly, for a density interaction at the $x_{0}=1$ boundary, it turns out that the ultra-strong boundary bosonic condensation is not just due to the interaction-NHSE interplay, but also relies crucially on SSH topology. This is evident from the 2-boson band structure plot in Fig. 2(b), where each eigenenergy $E$ is colored according to the overlap $\psi_{11}=\langle(1,1)|\psi\rangle$ of its corresponding eigenstate $|\psi\rangle$ with $|(1,1)\rangle=(b_{1}^{\dagger})^{2}|0\rangle$. Computed for our Hamiltonian $H$ [Eq. 1] at strong non-Hermitian hopping asymmetry $r=4$, it features 5 bands, with bands 2 and 4 from the hybridization of bulk and topological bands. This follows from the single-boson band structure [78, 38, 79, 80, 81, 82] with two symmetrically gapped bulk bands separated by in-gap topological zero modes. The effect of the $U$ density interaction is to induce high $\psi_{11}$ overlap (red) successively from bands 3 to 5, as $U$ is increased.
Figure 4: Persistence of ultrastrong boundary condensation $\bar{P}_{11}$ from interaction-induced caging, even for initial state (4,4) outside a $x_{0}=3$ cage. (a) Snapshots of the time-smoothed ($\Delta t=2$) two-boson correlation probability $\bar{P}_{11}$ (Eq. 6) for (a1) non-interacting and (a2) interacting cases. The bosons still gradually penetrate the L-shaped cage due to the NHSE towards (1,1), and remains trapped in it after that ($t=3,8$). (b1) Eventually, $\bar{P}_{11}(t)$ is ultra-enhanced only for the non-Hermitian interacting case (red), with (b2) negligible enhancement of boundary localization by $U$ in the single-boson case, similar to (b1,b2) of Fig. 3. (c) Time-averaged two-boson probability $\bar{P}_{\text{cage}}$ within the $x_{0}\times x_{0}$ cage as a function of hopping asymmetry $r$. For initial states $(2,2)$ and $(4,4)$ with coincident bosons, the $U=1$ cases exhibit much larger $\bar{P}_{\text{cage}}$ due to interaction-induced caging. But for $(2,4)$, non-coincident initial bosons do not interact appreciably and the effect of interactions is negligible.
@@
To illustrate, with the initial state (1,1), suppressed $\bar{P}_{11}$at $U\approx 3$ corresponds to high $\psi_{11}$ overlap with band 4 [circled in black in Figs. 2(a,b)]. This is because band 4 contains one bulk and one topological boson, which is not consistent with maintaining both bosons at $x_{0}=1$. As such, the same interaction strength of $U\approx 3$ leads to enhanced $\bar{P}_{11}$ (dashed black) for the initial state (1,4), where only one boson overlaps with the topological boundary mode. Likewise, $U\approx 1$ or $U\approx 5$ favors the $\bar{P}_{11}$ bosonic condensation for the bulk initial state (4,4), as circled in light blue, since that is when the $\psi_{11}$ overlap with the bulk bands (3 and 5) is strongest.

### v3

# Non-Hermitian ultra-strong clustering through interaction-induced caging

@@ Abstract
We uncover a new mechanism whereby the triple interplay of non-Hermitian pumping, bosonic interactions and nontrivial band topology leads to strong boson clustering. The extent of boson clustering goes beyond what is naively expected from the interaction-induced trapping of non-Hermitian pumped states and is based on an emergent caging mechanism that topological boundary modes can further enhance. Beyond our minimal model with two bosons, this caging remains applicable for generic many-boson systems subject to a broad range of density interactions and non-Hermitian hopping asymmetry. Our novel mechanism for particle clustering would inspire fundamental shifts in our comprehension of many-body non-Hermitian dynamics and open new avenues for controlling and manipulating bosons.
Keywords: Non-Hermitian dynamics, Bosonic interaction, Topological states, Non-Hermitian skin effect
I Introduction
There has been growing excitement in interplay between many-body interactions and the non-Hermitian skin effect (NHSE) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. This stems from the non-locality and sensitivity of the NHSE [2, 11], which profoundly impacts particle localization [3, 4, 5, 6] and localization dynamics [7, 8, 10], which leads to unexpected consequences in the presence of quantum many-body interactions. These include the emergence of a many-fermion Fermi skin [1, 6], exotic spin liquids from environmental couplings [3], spectral symmetry breaking [9], unconventional quantum dynamics in non-Hermitian baths [8, 10, 16] and the non-Hermitian Mott skin effect [15].
In this work, we uncovered a new mechanism whereby bosons can be made to condense very strongly through the triple interplay of non-Hermitian pumping, density interactions and non-trivial band topology. Due to the emergent non-locality of both the NHSE [2, 11] and the density interactions, the observed phenomenon contrasts starkly with the conventional behavior of interacting bosons, such as photons, which repel each other when they interact i.e. photon blockade [17, 18, 19, 20]. In particular, we show that ultra-strong clustering, and not just boundary state accumulation, arises from an emergent caging mechanism that requires both non-Hermitian pumping and density interactions.
II Methods
To demonstrate how non-Hermitian pumping, topology and bosonic interactions can interplay to cause unexpectedly strong bosonic clustering, we consider a minimal interacting 1D bosonic lattice model that contains topological edge modes at its boundary sites:
@@ II Methods
which is a Su-Schrieffer-Heeger (SSH) lattice [21] with asymmetric intercell ($t_{L}$ and $t_{R}$) hoppings and unit intracell hoppings, equipped with a density interaction $\frac{1}{2}Un_{x_{0}}^{2}$ at site $x_{0}$, $U\geq 0$. Here $b_{x}$ ($b_{x}^{\dagger}$) is the bosonic annihilation (creation) operator at sites $x=1,2,\ldots,2L$, with $n_{x_{0}}=b_{x_{0}}^{\dagger}b_{x_{0}}$ the boson number operator. Non-Hermitian pumping [22, 23, 24, 25, 26, 27, 28, 29, 11, 30, 31] occurs towards the left if $r=t_{L}/t_{R}>1$, accumulating the bosons towards the left boundary. However, when multiple indistinguishable bosons exist, this accumulation would also be affected by the repulsive ($U>0$) interaction at site $x_{0}$, with strength depending on bosonic occupancy. In this work, we have chosen the interaction $\frac{1}{2}Un_{x_{0}}^{2}$ that mimics the simplest possible nonlinearity in the mean-field limit, different from the usual Bose-Hubbard interaction [32, 17, 33, 18, 19] by a local density shift [34, 35, 36, 37, 38].
Figure 1: Two-boson correlation dynamics and associated spatial density distributions in a boundary-interacting system. The dramatic enhancement of bosonic clustering at site $x=1$ is captured by $P_{11}$ [Eq. 5], which measures the probability of finding both bosons at this site. This correlation is significantly stronger (red curves in panel e) when both the non-Hermitian skin effect (NHSE, $r>0$) and boundary interaction ($U>0$) are present, demonstrating their synergistic effect on boson trapping. This enhanced clustering is studied for different initial states $(1,1)$, $(4,4)$, and $(1,4)$ evolving under $H$ [Eq. 1]. Supporting this observation, panels (a-d) show the spatial density evolution $\rho_{x}(t)$ [Eq. 3] over time $t\in[0,10]$, with corresponding time-averaged profiles $\bar{\rho}_{i}$ [Eq. 4] plotted above. The density distributions contrast four scenarios: Hermitian ($r=1$) without (a) and with (b) interaction, and non-Hermitian ($r=4$) without (c) and with (d) interaction. Notably, the interaction term $U$, despite being typically repulsive, acts as an effective trap at $x_{0}=1$ during non-equilibrium evolution when combined with leftward NHSE, as evidenced by the bright regions in the heatmaps. Parameters: $t_{0}=3$, and for $r=4$, $t_{L}=1.6$ and $t_{R}=0.4$.
@@
where $v_{xx^{\prime}}(t)=v_{x^{\prime}x}(t)$ is the amplitude of the basis state $|(x,x^{\prime})\rangle=b_{x}^{\dagger}b_{x^{\prime}}^{\dagger}|0\rangle$ containing indistinguishable bosons at sites $x,x^{\prime}$.
We define the 2-boson density at site $x=1,2,\ldots,2L$ by
@@
with $0\leq\rho_{x}(t)\leq 2$. Associated with it is the time-averaged spatial density
@@
where $T$ is the simulation duration.
While $\rho_{x}(t)$ and $\bar{\rho}_{x}$ reveal where the bosons localize on the lattice, a high value of $\rho_{x}(t)$ or $\bar{\rho}_{x}$ can physically arise either due to strong single-boson localization at site $x$, or moderate double-boson clustering at the same site. To quantify this important distinction, we also examine the two-boson correlation probability
@@
which represents the probability of observing a boson at sites $x$ and $x^{\prime}$ at the same time $t$. To reveal underlying trends in the evolution of the two-boson correlation, we also define the time-smoothed correlation
@@
III Results
III.1 Ultra-strong non-Hermitian Bosonic clustering
We first consider scenarios where the density interaction is at the leftmost (boundary) site $x_{0}=1$. Figs. 1(a-d) showcase the evolution of the dynamical density $\rho_{x}(t)$ [Eq. 3] and its time-average $\bar{\rho}_{x}$ [Eq. 4] of three illustrative initial 2-boson state configurations.
@@ III.1 Ultra-strong non-Hermitian Bosonic clustering
For the initial state $(1,1)$ (Top Row) where both bosons are already on $x_{0}=1$, the density interaction indeed repels the bosons in the Hermitian limit, as evidenced in the suppressed $\bar{\rho}_{x}$ density at $x=x_{0}=1$ across the Hermitian interacting vs. non-interacting cases (orange vs. blue). Somewhat expected, this suppression vanishes when the NHSE counteracts the repulsion by pumping the bosons towards $x_{0}=1$ (green and red). However, for the initial state $(4,4)$ (Center Row) where both bosons are initially far from $x_{0}=1$, the boson density oscillates and spreads out, failing to accumulate appreciably at $x_{0}=1$ except when both the non-Hermiticity and interaction are present (red). Indeed, in this non-equilibrium scenario, the $U>0$ density term behaves more like a local potential well that traps the $\rho_{x}(t)$ at $x=x_{0}$, rather than repulsion. Qualitatively similar behavior is observed in the density evolution for initial state $(1,4)$ (Bottom Row), when one of the bosons is initially already at $x_{0}=1$ and is as such unaffected by the NHSE.
@@
What is striking, however, is the ultra-strong two-boson clustering at boundary site $x_{0}=1$ when non-Hermiticity ($r>1$) and the density interaction ($U>0$) are both present. This is revealed in the dynamical $P_{11}(t)$ [Eq. 5] plots in Fig. 1(e), which shows the probability of both bosons condensing at $x=1$. From the Top Row, when both bosons are prepared at $x=1$ (initial state =(1,1)), they can only remain there if the interaction and non-Hermiticity are simultaneously present (red). But most surprising is what happens when at least one boson is initially away from $x=1$ (Center and Bottom Rows): we observe overwhelmingly higher $P_{11}$ only when non-Hermiticity and interaction simultaneously interplay (red), compared to having either on their own (yellow, green). This suggests that trapping both bosons and suppressing photon blockade [18, 19, 17] hinges on an emergent consequence of this interplay, a fact not evident from the density evolution alone [Fig. 1(a-d)].
@@
To quantitatively characterize this two-boson clustering, i.e., $P_{11}(t)$ , we plot in Fig. 2(a) the time-averaged clustering probability $\bar{P}_{11}=\bar{P}_{11}(T)$ with $\Delta t=T$ [Eq. 6], which is the correlation $P_{11}$ smoothed over the entire simulation duration $T$. The heatmap of $\bar{P}_{11}$ in the interaction strength vs. non-Hermiticity $U$-$r$ plane [Fig. 2(a)] reveals significantly enhanced $\bar{P}_{11}$ (colored) for certain $(U,r)$ regions, compared to the $r=1$ Hermitian limit where $\bar{P}_{11}$ essentially vanishes (black). While $\bar{P}_{11}$ generally increases with $r$, optimal clustering $\bar{P}_{11}$ occurs at windows of $U$ that are highly dependent on the initial state, as will be explained below.
@@
Figure 2: Extent of two-boson clustering due to a boundary density interaction at $x_{0}=1$ [Eq. 1], and its correspondence with the boundary clustering of 2-boson spectral bands. (a) Time-averaged two-boson clustering probability $\bar{P}_{11}$ [Eq. 6] at site $x=1$, in the parameter space of non-Hermitian hopping asymmetry $r$ and density interaction strength $U$. (b) Corresponding two-boson spectra (purely real) at $r=4$ ($t_{L}=1.6$ and $t_{R}=0.4),t_{0}=3$, which features five bands, with bands 2 and 4 being containing a pair of interaction-hybridized bulk and topological bosons. As $U$ increases, it creates a group of eigenstates at $E\approx U$ that exhibits strong clustering $\psi_{11}$ (red) at $x=x_{0}=1$. Strong $\psi_{11}$ clustering in the hybrid topological band 4 leads to suppressed $\bar{P}_{11}$ for the boundary-localized initial state $(1,1)$ (circled in dashed black), but enhanced $\bar{P}_{11}$ clustering for the initial state $(1,4)$ (solid black). By contrast, strong $\psi_{11}$ clustering in the bulk bands 3 and 5 (circled in light blue) corresponds to enhanced $\bar{P}_{11}$ for the bulk initial state $(4,4)$.
III.2 Topological origin of ultra-strong bosonic clustering
[…]
@@ III.2 Topological origin of ultra-strong bosonic clustering
To illustrate, with the initial state (1,1), suppressed $\bar{P}_{11}$at $U\approx 3$ corresponds to high $\psi_{11}$ overlap with band 4 [circled in black in Figs. 2(a,b)]. This is because band 4 contains one bulk and one topological boson, which is not consistent with maintaining both bosons at $x_{0}=1$. As such, the same interaction strength of $U\approx 3$ leads to enhanced $\bar{P}_{11}$ (dashed black) for the initial state (1,4), where only one boson overlaps with the topological boundary mode. Likewise, $U\approx 1$ or $U\approx 5$ favors the $\bar{P}_{11}$ bosonic clustering for the bulk initial state (4,4), as circled in light blue, since that is when the $\psi_{11}$ overlap with the bulk bands (3 and 5) is strongest.
III.3 Two-boson interactions as caging mechanisms
Figure 3: Ultrastrong boundary clustering $\bar{P}_{11}$ from interaction-induced caging, for initial state (2,2) within a $x_{0}=3$ cage for a 5 unit cell chain. (a) Snapshots of the time-smoothed ($\Delta t=2$) two-boson correlation probability $\bar{P}_{11}$ [Eq. 6] for (a1) non-interacting and (a2) interacting cases. The L-shaped cage in (a2) traps the bosons tightly. (b1) Evolution of probability of both bosons at the left boundary, with greatly enhanced $\bar{P}_{11}(t)$ only for the non-Hermitian interacting case (red). (b2) When only one boson is present, the boundary density $\bar{\rho}_{1}(t)$ remains low regardless of whether a $U$ barrier is present, showcasing that the caging mechanism is an interaction effect.
Figure 4: Persistence of ultrastrong boundary clustering $\bar{P}_{11}$ from interaction-induced caging, even for initial state (4,4) outside a $x_{0}=3$ cage. (a) Snapshots of the time-smoothed ($\Delta t=2$) two-boson correlation probability $\bar{P}_{11}$ (Eq. 6) for (a1) non-interacting and (a2) interacting cases. The bosons still gradually penetrate the L-shaped cage due to the NHSE towards (1,1), and remain trapped in it after that ($t=3,8$). (b1) Eventually, $\bar{P}_{11}(t)$ is ultra-enhanced only for the non-Hermitian interacting case (red), with (b2) negligible enhancement of boundary localization by $U$ in the single-boson case, similar to (b1,b2) of Fig. 3. (c) Time-averaged two-boson probability $\bar{P}_{\text{cage}}$ within the $x_{0}\times x_{0}$ cage as a function of hopping asymmetry $r$. For initial states $(2,2)$ and $(4,4)$ with coincident bosons, the $U=1$ cases exhibit much larger $\bar{P}_{\text{cage}}$ due to interaction-induced caging. But for $(2,4)$, non-coincident initial bosons do not interact appreciably and the effect of interactions is negligible.